40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7877     Entry time: Mon Jan 7 00:08:16 2013
Author: Jenne 
Type: HowTo 
Category: Alignment 
Subject: In-vac plan 

List of things to do, in order:

* Remove BS heavy door.  Steve, please remove the BS door as soon as you have enough people to do so.  I will be a little late, since I have a dentist appointment, but please don't wait for me.  Jamie and Manasa can help you.  Put on a light door.

* Remove MC light doors, make aluminum foil tube (not light access connector, yet).

* Open laser shutter, lock PMC. (Required slight tweaking of input steering.) Confirm power level into vacuum <100mW.

* Lock MC and check spot positions of MC (quickly.  this shouldn't take all day, hopefully).

-------------------------------  End of work for Monday.  See following elog ------------------------------------------------

 

* Move TT1 to be as close as possible to the location indicated on the diagram, then align it. 

         * Make sure beam out of Faraday is hitting the center of the optic.

         * Make sure beam reflected off of TT1 hits center of PZT2.  Only use actuators for the final alignment, then confirm that they aren't close to the edge of their ranges.

         * Lock down TT1 with dog clamps.

* Put light access connector on MC.

* Swap PZT2 out with TT2.  Should be at correct spot, according to diagram, and beam should be hitting center of optic.  Alignment only to the ~few degree point here.

* Re-level BS table.

* Fix oplevs that need fixing.  (Manasa should have the plan on one of the diagrams).

* Put target on PRM cage.

* Align TT2 so that beam goes through PRM target.

* Open ITMX heavy door. (Probably Tuesday morning).

* Place freestanding target in front of PR2.  Ensure TT2 is aligned to go through PRM target, and hit center of PR2.  Again, save actuators for fine-tuning.

At this point, I think we should (temporarily) install one of the G&H mirrors as a flat mirror facing the PRM, and see if we can lock that cavity using REFL.  We will want to have already created a model for this case, to compare our observations to.  Or we could align the full PRMI, and try to lock that in air.

 

 

ELOG V3.1.3-